1. What types of items should be separately stated on Schedules K and K-1?

2.The partnership has ordinary income of $100,000before the payment of a guaranteed payment of $30,000 to Partner A. Partners A and B are 50-50 partners. What is the total income to Partner A and B for the year?

3.If a partnership has organizational expenses of $53,000, What is the total amortization expenses that can be deducted by the partnership if it begins business on November 10?
Answer #1

1. Schedule K and K1 is used to report items that are taxed separately for the partner such as share of income, profits ,deductions ,credits etc. These are items that are specially treated on a partners tax return.

2. Guaranteed payment to A = 30,000

Distributive share = 100,000 - 30,000 = 70,000

Distributive share to A = 70,000 * 1/2 = 35,000

Distributive share to B= 70,000 * 1/2 = 35,000

Total income for A = 30,000 + 35,000 = 65,000

Total income for B = 35,000

3. For organizational expenses exceeding $50,000, a deduction of $5000 can be made and for the remaining $48,000 the same is amortized over a 180 month period on a ratable basis.